OpenStudy (anonymous):

What are the solutions for x in the proportion 5x-6 2x ----- = --- x 2 x = −12 and x = −8 x = 12 and x = 8 x = −2 and x = −3 x = 2 and x = 3

OpenStudy (anonymous):

5x-6=x^2 x^2-5x+6=0 (x-2)(x-30=0 x=2,3

OpenStudy (hba):

the easy way to do it is to put the answer options in the question

OpenStudy (anonymous):

Thanks :)

OpenStudy (hba):